Determine the centroid (x, y) of the shaded area. Present your results in a table. 9 in 6 in
Answer #1


Sol - با 45 in GID لو سے - - ۹:۲ ) - 6n 4 (6) = 2 54 \ 37 4(6) 54 31] Ağ in? A Area, in Component +54 + 162 Triangle ļ (9) (6)= 27 +108 + 108 (6 = 36 Square +243 + 162 4.5 Rectangle 6x9 = 54 +108ř= E AX a 549.19 181.27 3 in E 360 2 - EA 181.27 ain Centroid = (3, 2) inch